OpenStudy (anonymous):

x-2y=-4 y=-3x+9 solve by substitution

OpenStudy (anonymous):

x = 2y - 4 From 1st. There fore, y = -3(2y - 4) + 9 7y = 21 y = 3 x = 6-4=2